OpenStudy (anonymous):

Which expressions are equivalent to 3(x + 3y + 2x − y )? 3(3x + 2y) and 9x + 6y 3(3x + 4y) and 9x + 12y 3(x + 4y) and 3x + 12y 3(x + 2y) and 3x + 6y

OpenStudy (anonymous):

@Kim21 I think its the last one? not sure.

OpenStudy (anonymous):

well first of all you would have to solve for x and then figure out the answer to 3(x + 3y + 2x − y ) in order to get your answer ;)

OpenStudy (anonymous):

so do you know how to do that?

OpenStudy (anonymous):

am going to do this using substitution and then by elimination. SUBSTITUTION METHOD Let's solve the first one in terms of y. y = 3-x. Plug 3-x where you see y in the second equation. 2x - 3(3-x) = 16 2x - 9 +3x = 16 5x -9 = 16 5x = 25 x = 5 Plug x=5 back into the other equation. 2(5) - 3y = 16 10 - 3y = 16 -3y = 6 y = -2 So our solution is (5,-2) or x=5y=-2

OpenStudy (anonymous):

now solve each of the choices that they give you until you get one with the same answer :)